OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) show all work! distribute and solve for x

OpenStudy (anonymous):

help! please

OpenStudy (callisto):

5(2x - 7) = 33 - 3(6x + 4) For left side, 5(2x-7) = 5(2x) - 5(7) = 10x - 35 For right side, 33 - 3(6x + 4) = 33 - 3(6x) -3(4) = Can you simplify it?

OpenStudy (anonymous):

33-18x-12

OpenStudy (anonymous):

?

OpenStudy (callisto):

33-18x-12 = (33 - 12) -18x =Can you simplify it?

OpenStudy (callisto):

It is not yet solved.. You're still on the distributing process.